SPECTROSCOPY OF THE A^{3}Π − X^{3}Σ^{−} TRANSITIONS OF NH-He AND NH-H_{2}

\underline{GALINA KERENSKAYA}, MICHAEL C.HEAVEN, Department of Chemistry, Emory University, Atlanta, GA 30322.

Complex features of NH-He were observed in conjunction with the A-X 1-0 and 0-0 transitions. Bands appear to be clustered around the lowest rotational line of the monomer. The energy range spanned by the sharp structure is only 6.5 cm^{−1}. The lowest energy feature is red shifted from the monomer by about 3 cm^{−1}.

A bound-free continuum spectrum was observed for NH-H_{2} in the region of the A-X 0-0 transition. Excitation to the A state is speculated to result in rapid NH-H_{2} -> NH_{2} + H reaction. The ground state dissociation energy of D′′_{0} = 32.4cm^{−1} was deduced from the spectrum.

Theoretical analyses of the NH-He and NH-H_{2} A-X bands are in progress.